OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

where (x1,y1) is a point given to you m= slope

OpenStudy (anonymous):

3x + y = 3